B) A baseball of mass 0.15 kg thrown at 100 mph has a momentum of 6.7 kg x m/s. if the uncertainty in measuring this momentum is 1.0 x 10-7 of the momentum, calculate the uncertainty in the baseballs position.
Answer #1

Answer B -

Mass of Baseball = 0.15 kg

Speed = 100 mph or 44.704 m/s [1 mph = 0.44704 m/s]

momentum = 6.7 kg.ms-1

uncertainty in measuring momentum = 1.0 * 10-7

uncertainty in position = ?

We know that,

\Deltap * \Delta x = h/4\pi ----------------------A

where, \Delta p = uncertainty in momentum

\Deltax = uncertainty in position

h = plank's constant (6.62607004 × 10-34 m2 kg / s)

\pi = 3.14

Also,

\Deltap = uncertainity in measurement * p

\Deltap = 1.0 * 10-7 * 6.7 kg.ms-1

\Deltap = 6.7 * 10-7 kg.ms-1

Now put all the values in A

\Deltap * \Delta x = h/4\pi

6.7 * 10-7 kg.ms-1 * \Delta x = 6.62607004 × 10-34 m2 kg/s / 4 * 3.14

\Deltax = 6.62607004 × 10-34 m2 kg/s / (4 * 3.14 * 6.7 * 10-7 kg.ms-1)

\Deltax = 7.87 * 10-29 m [Answer]